OpenStudy (anonymous):

write the equation of the line in point slope form: through (-1, 4) and (-1, 7)

OpenStudy (anonymous):

use the slope formula y2-y1/x2-x2

OpenStudy (anonymous):

I worked it out and the slope is undefined... so how do I put it into point slope form? (y-y1=m(x-x1)

OpenStudy (anonymous):

yes it means the line is vertical so all you write is x=-1 no matter what y values would be x will always equal -1

OpenStudy (anonymous):

but what do you substitute for m in the equation of point slope form because m is undefined!!! I need this answer!!!

OpenStudy (anonymous):

help me plz I will give a medal!

OpenStudy (anonymous):

nothing the equatio is x=-1 which is a vertical line

OpenStudy (anonymous):

so the equation in point slope form would be: y+1= (x+1) right???

OpenStudy (anonymous):

no forget about the slope. because the slope is undefined the equation will only be X=-1. if you graph the equation you only get a vertical line

OpenStudy (anonymous):

i know that i get a vertical line but i still don't get how to put it in that form... never mind

OpenStudy (anonymous):

i gave you a medal

OpenStudy (anonymous):

it wont go to the for of mx+b this is special straight line.

OpenStudy (anonymous):

The line is x = -1.

OpenStudy (anonymous):

i meant the form y2- y1= m (x-x1)

OpenStudy (anonymous):



OpenStudy (anonymous):

do see this drawing that how the line looks like

OpenStudy (anonymous):

i know what it looks like, i need the point slope form of the equation! i get that in slope intercept it would be x=-1, but i just don't get how to do it for point slope form! can you write the equation in point slope form for me?

OpenStudy (anonymous):

It can't be written in the point slope form.

OpenStudy (anonymous):

so... I leave the answer blank?

OpenStudy (anonymous):

undifined slope

OpenStudy (anonymous):

Explain and draw the line and show that the equation is x = -1. Slope is undefined....etc. etc.! You have a lot to write than leave blank.

OpenStudy (anonymous):

my assignment was to write the equation of the line in point slope form, slope intercept form, and standard form... pretend this is the assignment, and write those answers for me as the question should be answered

OpenStudy (anonymous):

no drawing needed, just equation

OpenStudy (anonymous):

Standard form will be x=-1

OpenStudy (anonymous):

what about point slope and slope intercept! im kinda getting annoyed... no offense...

OpenStudy (anonymous):

hhhhhhhhheeeeeeeeeeeeellllllllllllllllllllllllllllpppppppppppppp!!!

OpenStudy (anonymous):

How many ways should I say it is undefined in those forms?

OpenStudy (anonymous):

what does your question says

OpenStudy (anonymous):

he/she cant understand

OpenStudy (anonymous):

keeping in mind that the 2 points given were (-1, 4) and (-1, 7) and the slope is undefined, how do i put rewrite that info into the equation: y-y1= m (x- x1) m is the slope, y1 is 7, and x1 is -1

OpenStudy (anonymous):

We know. m is undefined. so the whole thing is undefined.

OpenStudy (anonymous):

no no

OpenStudy (anonymous):

so i should write as my answer: undefined ???

OpenStudy (anonymous):

forget about rewriting the equation just go with X=-1 thats the answer it does not apply to this type of graph

OpenStudy (anonymous):

yes m- undifined

OpenStudy (anonymous):

okay i'll just ask my mom never mind gotta go bye

OpenStudy (anonymous):

have fun